OpenStudy (anonymous):

determinate k so that the polynomial can be writen q(x) * (x+2) and determinate q(x) x^2 + 8x + k

Parth (parthkohli):

\[q(x) \times (x + 2) = x^2 + 8x + k\implies q(x) = {x^2 + 8x + k \over x+2} \]

Parth (parthkohli):

Is that a good hint?

OpenStudy (anonymous):

then long devision?

Parth (parthkohli):

Yeah. I am still not sure what the question means yet.

Parth (parthkohli):

Is \(q(x) = x^2 + 8x + k\)?

OpenStudy (anonymous):

yeah my english is not good

OpenStudy (anonymous):

answers are k=12 and q(x) = x+ 6

OpenStudy (turingtest):

\[Q(x)(x+2)=(x+a)(x+2)=x^2+8x+k\]we know the middle term comes from\[a+2=8\implies a=6\]

Parth (parthkohli):

Still trying to decipher. Let's see what Turing says.

OpenStudy (turingtest):

once you know that you can easily foil out\[(x+6)(x+2)\]and find that \(k=12\)

OpenStudy (anonymous):

using long devision the rest is 12 and the function is x + 6 maybe thats what they want?

OpenStudy (anonymous):

Thanks everyone !!!

OpenStudy (turingtest):

welcome!

Parth (parthkohli):

Oh, I gotcha, Turing!

OpenStudy (turingtest):

cool, let me know if you want further clarification kalemale